# Continuous Urban Change Detection from Satellite Image Time Series with Temporal Feature Refinement and Multi-Task Integration

### v1

@@ Abstract
Urbanization advances at unprecedented rates, resulting in negative effects on the environment and human well-being. Remote sensing has the potential to mitigate these effects by supporting sustainable development strategies with accurate information on urban growth. Deep learning-based methods have achieved promising urban change detection results from optical satellite image pairs using convolutional neural networks (ConvNets), transformers, and a multi-task learning setup. However, transformers have not been leveraged for urban change detection with multi-temporal data, i.e., $>2$ images, and multi-task learning methods lack integration approaches that combine change and segmentation outputs. To fill this research gap, we propose a continuous urban change detection method that identifies changes in each consecutive image pair of a satellite image time series. Specifically, we propose a temporal feature refinement (TFR) module that utilizes self-attention to improve ConvNet-based multi-temporal building representations. Furthermore, we propose a multi-task integration (MTI) module that utilizes Markov networks to find an optimal building map time series based on segmentation and dense change outputs. The proposed method effectively identifies urban changes based on high-resolution satellite image time series acquired by the PlanetScope constellation (F1 score 0.551) and Gaofen-2 (F1 score 0.440). Moreover, our experiments on two challenging datasets demonstrate the effectiveness of the proposed method compared to bi-temporal and multi-temporal urban change detection and segmentation methods. Code is available on GitHub: https://github.com/SebastianHafner/ContUrbanCD.
@@ I Introduction
In recent years, deep learning has been continuously replacing traditional change detection methods [8, 9]. Specifically, deep convolutional neural networks (ConvNets) have been used extensively for change detection in bi-temporal optical satellite image pairs (see Figure 1(a)). The simplest way of adapting common ConvNets such as U-Net [10] or HRNet [11] for change detection is with an input-level fusion (or early fusion [12]) strategy, referring to the concatenation of image pairs before passing them to a ConvNet. Contrarily, late fusion strategies typically process images separately in a Siamese network consisting of two ConvNets with shared weights. Extracted bi-temporal features are then fused using concatenation or absolute differencing [12, 13]. Since Siamese networks are generally considered preferable to input-level fusion strategies, multiple studies developed modules that are incorporated into Siamese network architectures to improve feature representations [14, 15, 16]. For very high-resolution (VHR) imagery, Chen et al. [17, 18] proposed to refine features extracted by ConvNets using a transformer-based module that alleviates the limited long-range context modeling capability of convolutions with self-attention. Since then, self-attention has become a popular mechanism for capturing long-range contextual information in VHR change detection [19, 16, 20, 21].
@@
In recent years, high-resolution (i.e., 1 – 10 m) satellite image time series have become increasingly available [22]. Those data have enabled a shift from detecting land cover changes in image pairs acquired years apart to continuous annual and sub-annual change detection [23]. In contrast, urban change detection methods are predominantly designed for bi-temporal change detection from image pairs acquired multiple years apart [24]. However, considering the unprecedented rate of global urbanization [25], it is essential to develop a new suite of methods that detect urban changes continuously. While continuous change detection can be achieved by applying a bi-temporal model to consecutive image pairs in satellite image time series (see Figure 1(c)), this approach fails to exploit multi-temporal, i.e., $>$ 2 images, information. Furthermore, recent studies [26, 27] have demonstrated the effectiveness of multi-temporal change detection models which predict changes between the first and last images of a satellite image time series (see Figure 1(b)). For example, multi-temporal information can help to reduce commission errors from registration errors, illumination differences, or other types of change unrelated to the problem of interest [27]. In addition, it can mitigate the effect of cloud artifacts in single images [26]. Existing multi-temporal change detection methods employ either recurrent layers, such as long short-term memory (LSTM) layers [28], or 3D convolutional layers to model temporal information [26, 27, 29, 30]. While these layers effectively model short-range temporal dependencies in time series data, the self-attention mechanism can explicitly model temporal dependencies across all timestamps of a time series [31]. However, self-attention is predominately used in urban change detection to model long-range dependencies in space [17, 18, 19, 16]. In contrast, this research aims to use self-attention to model multi-temporal information extracted from satellite image time series for urban change detection.
@@
Another promising avenue of research for change detection is multi-task learning [32], where a related semantic segmentation task is trained parallel to the change detection task using a shared feature representation. The change detection task is commonly combined with building segmentation [33, 27, 34, 35, 36] or land cover segmentation [37, 38, 39]. To that end, Siamese networks are extended with an additional decoder for the semantic segmentation task. The feature maps extracted by the encoder are then shared between the change decoder and the segmentation decoder. However, despite the attention multi-task learning has attracted in change detection, effective methods to integrate segmentation and change outputs have been largely unaddressed. For example, multi-task binary change detection studies consider building and change predictions independent outputs of the network [33, 27, 34, 35, 36]. On the other hand, semantic change detection studies, concerned with identifying not only whether changes occurred (i.e., binary change detection) but also what types of changes occurred, apply a simple masking approach for multi-task integration [37, 38, 39]. Specifically, land cover predictions and binary change predictions are integrated by masking land cover changes derived from the bi-temporal semantic predictions with the binary change prediction. Therefore, existing multi-task change detection studies do not account for inconsistencies between the segmentation and change predictions. Moreover, these approaches fail to exploit the complementary information produced by multi-task predictions.
@@
In this paper, we propose a continuous urban change detection method (see Figure 1(d)) and explore two research gaps in the current literature, namely (1) the use of the self-attention mechanism to model multi-temporal information in urban change detection and (2) the integration of segmentation and change predictions in multi-task learning setups. Specifically, we propose a new network architecture that relies on convolutions to extract multi-temporal building representations and employs self-attention to model temporal dependencies in feature space. We also propose a novel integration approach that determines the most likely state for each timestamp in a time series by considering multi-task predictions using Markov networks. The effectiveness of the proposed architecture and integration approach is demonstrated on two urban change detection datasets featuring high-resolution optical satellite image time series, namely the SpaceNet 7 dataset [40] and the Wuhan Urban Semantic Understanding (WUSU) dataset [41].
@@
  • We present a transformer-based feature refinement module that effectively models temporal information in satellite image time series, leading to improved building representations.
  • We introduce a new multi-task integration approach that represents segmentation and change outputs in Markov networks to find the most likely building map for each image in a time series using maximum a posteriori (MAP) estimation.
@@ II-A Bi-temporal change detection
II-B Change detection from satellite image time series
Few studies have developed deep learning methods for urban change detection from high-resolution satellite image time series. For example, Papadomanolaki et al. [26] proposed to incorporate LSTM networks into a U-Net model to leverage optical satellite image time series for change detection. Their L-UNet outperformed bi-temporal ConvNet-based methods on a bi-temporal dataset enriched with intermediate satellite images [26]. Others proposed an encoder-decoder LSTM model that is trained to rearrange temporally shuffled time series [30]. The core assumption of this unsupervised method is that the model fails to correctly rearrange shuffled data for changed pixels. Finally, Meshkini et al. [29] recently proposed a weakly supervised change detection method that employs 3D convolutional layers to capture spatial-temporal information in satellite image time series. Despite these promising results underlying the value of time series information, multi-temporal urban change detection methods may benefit from layers with better temporal modeling capabilities. Furthermore, most of these methods fail to detect urban changes continuously and are limited to the detection of changes between the first and the last image (i.e., multi-temporal change detection in Figure1(b)).
@@ II-C Multi-task learning
Apart from binary change detection, multi-task learning has been employed for semantic change detection [37, 38, 39] and 3D change detection [20]. Daudt et al. [37] proposed a learning strategy that integrates change detection and land cover mapping from bi-temporal image pairs. Specifically, they propose a network consisting of separate encoder-decoder sub-networks for land cover mapping and binary change detection, where the change detection decoder also receives features from the mapping encoder in the form of difference skip connections. ChangeMask [38] uses a similar setup but introduces a transformer applied to the bi-temporal feature maps to learn more robust representations for the change task. Furthermore, [39] optimized semantic features in ChangeMask by adding a multi-task join loss (i.e., ChangeMJ). Finally, Marsocci et al. [20] extended the bi-temporal image transformer (BIT) presented in [18] with a regression head to solve 2D and 3D change detection simultaneously using shared feature representations.
@@ III-A Overview
As illustrated in Figure 2, the main components of the proposed method are a ConvNet-based encoder, transformer-based temporal feature refinement (TFR) modules, a change feature (CF) module, two task-specific ConvNet-based decoders, and a Markovian multi-task (MTI) integration module. The following summarizes the urbanization monitoring process of the proposed method for a time series of satellite images:
@@
    Finally, the building and urban change outputs are combined using the MTI module. The module uses pixel-wise Markov networks to obtain optimal building states for the satellite image time series.
@@ III-B Satellite image time series encoding
It should be noted that for brevity, Figure 2 illustrates the proposed method for $s\in\{0,1,2\}$.
@@ III-C Temporal feature refinement
The TFR module, illustrated in Figure 3, creates temporally refined feature maps using the self-attention mechanism along the temporal dimension [31]. Specifically, the module takes as input a time series of feature maps at the same scale $s$ and reshapes this 4D tensor to a 3D tensor $\mathbf{T}^{s}\in\mathbb{R}^{T\times D^{s}\times P^{s}}$ by flattening the spatial dimensions $H^{s}$ and $W^{s}$. Consequently, $T$, $D$, and $P$ represent the temporal, feature embedding, and spatial dimensions, respectively.
After reshaping the feature map time series, self-attention is applied along the temporal dimension $T$ for each cell in the spatial dimension $P$. However, since the self-attention mechanism contains no recurrence, it is necessary to first inject information about the temporal position of the feature vectors in the time series. Specifically, temporal encodings having the same dimension as the feature vectors are generated based on sine and cosine functions of different frequencies [31]. These relative temporal encodings are then added to the feature vectors.
The tensor, enriched with relative temporal position information, is passed through two transformer encoder layers (see Figure 4). The key component of the transformer encoder layer is the multi-head attention block, which performs self-attention defined as follows:
@@
Figure 3: Illustration of the temporal feature refinement (TFR) module.
@@ III-D Change feature extraction
We refer to the scenario when all possible combinations of image pairs are considered as dense. However, we also investigate sparser settings such as adjacent, considering only changes between temporally adjacent images ($N_{\rm adjacent}=T-1$), and cyclic, adding the changes between the first and last image to the adjacent setting ($N_{\rm cyclic}=T$) (see Figures 5(a)5(c) for visualizations of these settings in Markov networks with $T=5$). We provide ablation results on the different edge settings in Table III. Finally, it should be noted that the CF module does not have any trainable parameters.
@@ III-G Multi-task integration
The core idea of the MTI module is to represent segmentation and change outputs in a pairwise Markov network. This subclass of Markov networks is associated with an undirected graph $G=(\mathcal{N},\mathcal{E})$ in which the nodes $\mathcal{N}$ correspond to random variables and the edges $\mathcal{E}$ represent pairwise relationships between the nodes (see [46]). For a given pixel, we construct a Markov network with $T$ nodes corresponding to the timestamps in an image time series. Specifically, all nodes in the network correspond to a binary variable representing the presence of buildings (i.e., $\mathcal{N}^{t}\in\{true,false\}$). We use state 1 to denote $true$, representing the presence of a building, and state 0 to denote $false$, representing the absence of a building. Adjacent nodes in the network are connected with $N-1$ edges, where we use $\mathcal{E}^{t\rightleftharpoons k}$ to denote an edge connecting timestamps $t$ and $k$. We refer to this Markov network structure as an adjacent network (see Figure 5(a)).
@@
To incorporate the change outputs for a specific pixel, we first include additional edges for the edge settings cyclic and dense. Specifically, for the cyclic case (Figures 5(b)), edge $\mathcal{E}^{t1\rightleftharpoons T}$, connecting the first node $\mathcal{N}^{t1}$ and the last node $\mathcal{N}^{T}$, is added. On the other hand, for the dense case (Figure 5(c)), all possible non-adjacent edges are added to the graph. Then, we define factors over the edges in the Markov network to add pairwise interactions of connected nodes. Specifically, we define pairwise potentials $\phi_{n}$ for each edge $\mathcal{E}^{n}$, connecting two nodes $\mathcal{N}^{t}$ and $\mathcal{N}^{k}$. Since all variables in the network are binary, each factor over an edge has four parameters. The change outputs are encoded as edge potentials for the four combinations of states, as follows:
@@
To define a global model from the local interactions defined in the parameterization of the Markov network, we take the product of the local factors and then normalize it. Once the distribution is defined, we perform a MAP query to find the most likely state assignment for each node in the graph. To that end, inference is performed using the belief propagation algorithm (see Algorithm 10.4 in [46]). Due to the absence of loops in the graph, belief propagation provides an exact solution. Finally, it should be noted that the MTI module does not contain any trainable parameters and is only deployed during inference.

### v2

@@ Abstract
Urbanization advances at unprecedented rates, resulting in negative effects on the environment and human well-being. Remote sensing has the potential to mitigate these effects by supporting sustainable development strategies with accurate information on urban growth. Deep learning-based methods have achieved promising urban change detection results from optical satellite image pairs using convolutional neural networks (ConvNets), transformers, and a multi-task learning setup. However, transformers have not been leveraged for urban change detection with multi-temporal data, i.e., $>2$ images, and multi-task learning methods lack integration approaches that combine change and segmentation outputs. To fill this research gap, we propose a continuous urban change detection method that identifies changes in each consecutive image pair of a satellite image time series (SITS). Specifically, we propose a temporal feature refinement (TFR) module that utilizes self-attention to improve ConvNet-based multi-temporal building representations. Furthermore, we propose a multi-task integration (MTI) module that utilizes Markov networks to find an optimal building map time series based on segmentation and dense change outputs. The proposed method effectively identifies urban changes based on high-resolution SITS acquired by the PlanetScope constellation (F1 score 0.551) and Gaofen-2 (F1 score 0.440). Moreover, our experiments on two challenging datasets demonstrate the effectiveness of the proposed method compared to bi-temporal and multi-temporal urban change detection and segmentation methods. Code is available on GitHub: https://github.com/SebastianHafner/ContUrbanCD.
@@ I Introduction
In recent years, deep learning has been continuously replacing traditional change detection methods [8, 9, 10, 11]. Specifically, deep convolutional neural networks (ConvNets) have been used extensively for change detection in bi-temporal optical satellite image pairs (see Fig. 1(a)). The simplest way of adapting common ConvNets such as U-Net [12] for change detection is with an input-level fusion (or early fusion [13]) strategy, referring to the concatenation of image pairs before passing them to a ConvNet. Contrarily, late fusion strategies typically process images separately in a Siamese network consisting of two ConvNets with shared weights. Extracted bi-temporal features are then fused using concatenation or absolute differencing [13, 14]. Since Siamese networks are generally considered preferable to input-level fusion strategies, multiple studies developed modules that are incorporated into Siamese network architectures to improve feature representations [15, 16, 17]. For example, Chen et al. [18, 19] proposed to refine features extracted by ConvNets from very high-resolution (VHR) imagery using a transformer-based module, alleviating the limited long-range context modeling capability of convolutions with self-attention. Since then, self-attention has become a popular mechanism for capturing long-range contextual information in VHR change detection [20, 17, 21, 22].
@@
In recent years, high-resolution (i.e., 1 – 10 m) satellite image time series (SITS) have become increasingly available [23]. Those data have enabled a shift from detecting land cover changes in image pairs acquired years apart to continuous annual and sub-annual change detection [24]. In contrast, urban change detection methods are predominantly designed for bi-temporal change detection from image pairs acquired multiple years apart [25]. However, considering the unprecedented rate of global urbanization [26], it is essential to develop a new suite of methods that detect urban changes continuously. While continuous change detection can be achieved by applying a bi-temporal model to consecutive image pairs in SITS (see Fig. 1(c)), this approach fails to exploit multi-temporal, i.e., $>$ 2 images, information. Furthermore, recent studies [27, 28] have demonstrated the effectiveness of multi-temporal change detection models which predict changes between the first and last images of a SITS (see Fig. 1(b)). For example, multi-temporal information can help to reduce commission errors from registration errors, illumination differences, or other types of change unrelated to the problem of interest [28]. In addition, it can mitigate the effect of cloud artifacts in single images [27]. Existing multi-temporal change detection methods employ either recurrent layers, such as long short-term memory (LSTM) layers [29], or 3D convolutional layers to model temporal information [27, 28, 30, 31]. While these layers effectively model short-range temporal dependencies in time series data, the self-attention mechanism can explicitly model temporal dependencies across all timestamps of a time series [32]. Thus, several recent segmentation methods for SITS employ the self-attention mechanism to explicitly model temporal dependencies across all timestamps of a time series [33, 34, 35]. However, the temporal models in these methods collapse the temporal dimension, resulting in a single output feature. Therefore, they do not facilitate continuous urban change detection which requires the full temporal information to produce change maps between each consecutive image pair in the SITS.
@@
Another promising avenue of research for change detection is multi-task learning [36], where a related semantic segmentation task is trained parallel to the change detection task using a shared feature representation. The change detection task is typically combined with building segmentation for urban change detection [37, 28, 38, 39, 40]. To that end, Siamese networks are extended with an additional decoder for the semantic segmentation task. The feature maps extracted by the encoder are then shared between the change decoder and the segmentation decoder. However, despite the attention multi-task learning has attracted in change detection, effective methods to integrate segmentation and change outputs have been largely unaddressed. For example, most multi-task urban change detection studies consider building and change predictions independent outputs of the network [37, 28, 38, 39, 40]. Therefore, these studies do not account for inconsistencies between the building segmentation and urban change predictions. Moreover, they fail to exploit the complementary information produced by multi-task predictions.
@@
In this paper, we propose a continuous urban change detection method (see Fig. 1(d)) and explore two research gaps in the current literature, namely (1) the use of the self-attention mechanism to model multi-temporal information in urban change detection and (2) the integration of segmentation and change predictions in multi-task learning setups. Specifically, we propose a new network architecture that relies on convolutions to extract multi-temporal building representations and employs self-attention to model temporal dependencies in feature space while preserving the temporal dimension. We also propose a novel integration approach that determines the optimal building segmentation for each image in a time series based on the multi-task network outputs. The effectiveness of the proposed architecture and integration approach is demonstrated on two urban change detection datasets featuring high-resolution optical SITS, namely the SpaceNet 7 dataset [41] and the Wuhan Urban Semantic Understanding (WUSU) dataset [42].
@@
  • We introduce a continuous urban change detection model that produces change outputs for each consecutive image pair in SITS while leveraging the full temporal dimension of the time series.
  • To enable continuous change detection, we present a transformer-based feature refinement module that effectively models temporal information in SITS. Importantly, our module preserves the temporal dimension of the representations, in contrast to existing temporal modules that collapse multi-temporal representations into a single one.
  • We propose a new multi-task integration approach that represents segmentation and change outputs in Markov networks to find the optimal building maps time series based on the network outputs.
@@ II-A Bi-temporal change detection
II-B Change detection and segmentation from time series data
Few studies have developed deep learning methods for urban change detection from high-resolution SITS. For example, Papadomanolaki et al. [27] proposed to incorporate LSTM networks into a U-Net model to leverage optical SITS for change detection. Their L-UNet outperformed bi-temporal ConvNet-based methods on a bi-temporal dataset enriched with intermediate satellite images [27]. Others proposed an encoder-decoder LSTM model that is trained to rearrange temporally shuffled time series [31]. The core assumption of this unsupervised method is that the model fails to correctly rearrange shuffled data for changed pixels. On the other hand, Meshkini et al. [30] proposed a weakly supervised change detection method that employs 3D convolutional layers to capture spatial-temporal information in SITS. Recently, He et al. [46] presented a deep learning method for time series land cover change detection. However, since their model only uses one-dimensional convolutions along the temporal dimension, it does not consider the spatial dimension, which is a limiting factor for high-resolution data.
Due to the limited number of change detection methods for SITS, we also expand this review to the semantic segmentation of SITS. Several recent semantic segmentation methods for SITS employed the self-attention mechanism for temporal modeling of multi-temporal features [33, 34, 35]. Garnot et al. [33] employed a lightweight-temporal attention encoder [47] for the temporal modeling of multi-temporal feature maps extracted using a shared ConvNet encoder. Similarly, Cai et al. [35] employed an attention bidirectional LSTM module for temporal modeling of ConvNet-based feature maps time series. The modules in both studies collapse the temporal dimension, resulting in a single feature map obtained using a ConvNet decoder. While Tarasiou et al. [34] used a vision transformer to learn feature representation from SITS, their model also outputs a single feature map for semantic segmentation.
In summary, existing multi-temporal change detection methods rely on recurrent and 3D convolutional layers for temporal modeling. While multi-temporal semantic segmentation methods frequently employ temporal attention-based modules, they collapse the temporal dimension similarly to multi-temporal change detection methods. While these methods can be adapted for change detection, collapsing the temporal dimension limits them to the detection of changes between the first and the last image of a SITS (i.e., multi-temporal change detection in Fig. 1(b)).
@@ II-C Multi-task learning
In general, these multi-task studies demonstrate that learning a segmentation task in parallel to the change detection task improves the latter. However, none of these studies investigate combining the change and segmentation network outputs to improve performance. Consequently, inconsistencies between the network outputs are also not accounted for.
@@ III-A Overview
As illustrated in Fig. 2, the main components of the proposed method are a ConvNet-based encoder, transformer-based temporal feature refinement (TFR) modules, a change feature (CF) module, two task-specific ConvNet-based decoders, and a Markovian multi-task (MTI) integration module. The following summarizes the urbanization monitoring process of the proposed method for a time series of satellite images:
@@
    Finally, the building and urban change outputs are combined using the MTI module. The module uses pixel-wise Markov networks to obtain optimal building states for the SITS.
@@ III-B Satellite image time series encoding
It should be noted that for brevity, Fig. 2 illustrates the proposed method for $s\in\{0,1,2\}$.
@@ III-C Temporal feature refinement
The TFR module, illustrated in Fig. 3, creates temporally refined feature maps using the self-attention mechanism along the temporal dimension [32]. Unlike the temporal modules in existing change detection and segmentation methods for SITS, our module preserves the temporal dimension.
The module takes as input a time series of feature maps at the same scale $s$ and reshapes this 4D tensor to a 3D tensor $\mathbf{T}^{s}\in\mathbb{R}^{T\times D^{s}\times P^{s}}$ by flattening the spatial dimensions $H^{s}$ and $W^{s}$. Consequently, $T$, $D$, and $P$ represent the temporal, feature embedding, and spatial dimensions, respectively. After reshaping the feature map time series, self-attention is applied along the temporal dimension $T$ for each cell in the spatial dimension $P$. However, since the self-attention mechanism contains no recurrence, it is necessary to first inject information about the temporal position of the feature vectors in the time series. Specifically, temporal encodings having the same dimension as the feature vectors are generated based on sine and cosine functions of different frequencies [32]. These relative temporal encodings are then added to the feature vectors.
The tensor, enriched with relative temporal position information, is passed through two transformer encoder layers (see Fig. 4). The key component of the transformer encoder layer is the multi-head attention block, which performs self-attention defined as follows:
@@
Figure 3: Illustration of the temporal feature refinement (TFR) module, preserving the temporal dimension of the input feature maps.
@@ III-D Change feature extraction
We refer to the scenario when all possible combinations of image pairs are considered as dense. However, we also investigate sparser settings such as adjacent, considering only changes between temporally adjacent images ($N_{\rm adjacent}=T-1$), and cyclic, adding the changes between the first and last image to the adjacent setting ($N_{\rm cyclic}=T$) (see Fig. 5(a)5(c) for visualizations of these settings in Markov networks with $T=5$). We provide ablation results on the different edge settings in Tables III and V. Finally, it should be noted that the CF module does not have any trainable parameters.
@@ III-G Multi-task integration
The core idea of the MTI module is to represent segmentation and change outputs in a pairwise Markov network. This subclass of Markov networks is associated with an undirected graph $G=(\mathcal{N},\mathcal{E})$ in which the nodes $\mathcal{N}$ correspond to random variables and the edges $\mathcal{E}$ represent pairwise relationships between the nodes (see [49]). For a given pixel, we construct a Markov network with $T$ nodes corresponding to the timestamps in an image time series. Specifically, all nodes in the network correspond to a binary variable representing the presence of buildings (i.e., $\mathcal{N}^{t}\in\{true,false\}$). We use state 1 to denote $true$, representing the presence of a building, and state 0 to denote $false$, representing the absence of a building. Adjacent nodes in the network are connected with $N-1$ edges, where we use $\mathcal{E}^{t\rightleftharpoons k}$ to denote an edge connecting timestamps $t$ and $k$. We refer to this Markov network structure as an adjacent network (see Fig. 5(a)).
@@
To incorporate the change outputs for a specific pixel, we first include additional edges for the edge settings cyclic and dense. Specifically, for the cyclic case (Fig. 5(b)), edge $\mathcal{E}^{t1\rightleftharpoons T}$, connecting the first node $\mathcal{N}^{t1}$ and the last node $\mathcal{N}^{T}$, is added. On the other hand, for the dense case (Fig. 5(c)), all possible non-adjacent edges are added to the graph. Then, we define factors over the edges in the Markov network to add pairwise interactions of connected nodes. Specifically, we define pairwise potentials $\phi_{n}$ for each edge $\mathcal{E}^{n}$, connecting two nodes $\mathcal{N}^{t}$ and $\mathcal{N}^{k}$. Since all variables in the network are binary, each factor over an edge has four parameters. The change outputs are encoded as edge potentials for the four combinations of states, as follows:
@@
To define a global model from the local interactions defined in the parameterization of the Markov network, we take the product of the local factors and then normalize it. Once the distribution is defined, we perform a maximum a posteriori query to find the optimal state assignment for each node in the graph. To that end, inference is performed using the belief propagation algorithm (see Algorithm 10.4 in [49]). Due to the absence of loops in the graph, belief propagation provides an exact solution. Finally, it should be noted that the MTI module does not contain any trainable parameters and is only deployed during inference.